# INACIA: Integrating Large Language Models in Brazilian Audit Courts: Opportunities and Challenges

### 19 Jan 2024 v2

@@ Abstract.
This paper introduces INACIA (Instrução Assistida com Inteligência Artificial), a groundbreaking system designed to integrate Large Language Models (LLMs) into the operational framework of Brazilian Federal Court of Accounts (TCU). The system automates various stages of case analysis, including basic information extraction, admissibility examination, Periculum in mora and Fumus boni iuris analyses, and recommendations generation. Through a series of experiments, we demonstrate INACIA’s potential in extracting relevant information from case documents, evaluating its legal plausibility, and formulating propositions for judicial decision-making. Utilizing a validation dataset alongside LLMs, our evaluation methodology presents an innovative approach to assessing system performance, correlating highly with human judgment. The results highlight INACIA’s proficiency in handling complex legal tasks, indicating its suitability for augmenting efficiency and judicial fairness within legal systems. The paper also discusses potential enhancements and future applications, positioning INACIA as a model for worldwide AI integration in legal domains.
copyright: nonejournalyear: 2023doi: XXXXXXX.XXXXXXXjournal: DGOVjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@ 1. Introduction
In this paper, we introduce a groundbreaking application of LLMs for administrative case processing and assisted decision-making developed jointly by the Brazilian Federal Court of Accounts (TCU, or Tribunal de Contas da União, in Portuguese) and experts at several Brazilian research organizations. INACIA (Instrução Assistida com Inteligência Artificial, in Portuguese or Instruction Assisted with Artificial Intelligence) is a system that suggests judication directions, which consist of a proposed way for the decision-making authority to address the case. Such directions are fundamentally the responsibility of TCU’s staff auditors. INACIA serves as an assistant in this process, providing analytical support. Its inputs are case filings; outputs are suggestions to the auditor to write the initial instruction [^1] The initial instruction is the document that contains the directions written by the auditor, including the main facts of the case, the legal basis, and the proposals.. INACIA uses a combination of LLMs and search engines to extract information from case documents and to reason through the extracted information.
@@
2. Background on Brazilian Audit Courts
@@ 2.2. Artificial Intelligence and TCU
Acknowledging the rapid evolution of AI and in line with its mission to foster innovation and enhance Public Administration, TCU initiated a series of studies in 2019 on Pre-commercial Procurement (PCP). Subsequently, in 2022, it launched its first “ETEC” (Encomenda Tecnológica, in Portuguese), a type of public procurement that requires Research and Development with technological risk, where the demanded solution extends beyond the current state of the art. Indeed, the envisaged solution is intended to expedite audit case processing by assisting the TCU auditor responsible for the case as effectively as a human specialist in such tasks would be capable of offering. After a competitive and meticulous selection process [^3] Available at https://tcu.gov.br/etec (content in Portuguese), the consortium of companies NeuralMind [^4] https://neuralmind.ai, a pioneer in solutions employing LLMs in Brazil, and TerraNova [^5] https://trnv.com.br/, specialized in jurimetrics, was chosen. INACIA represents the initial proof of concept for the desired outcome of the ETEC initiative. TCU aims to focus auditors’ expertise on the most challenging tasks of case processing while handing over the more routine tasks to the AI system. This strategy is well-aligned with the evidence on how best to incorporate technology for productivity gains and labor output augmentation (Acemoglu and Autor, 2011). The current expectation is that AI will change how TCU handles cases from a sequential, step-by-step approach to a more holistic approach in which filings are evaluated simultaneously and connected to external resources necessary for decision-making from the outset. Effectively, this system would eliminate redundancies and streamline case processing without requiring regulation change.
@@ 2.3. Audit Case Processing
TCU investigates (and eventually issues rulings) the use of federal funds in two ways: (1) when TCU independently audits federal agencies (or anyone using federal funds) or (2) when TCU receives complaints filed by individuals, private entities, or public agencies on the misuse of federal funds.
@@
While case proceedings may vary based on the initiation method, case processing is typically a time-intensive task primarily handled by TCU staff. Auditors must review case filings, confirm whether TCU should hear the case, search and append external evidence of wrongdoing (if necessary), check claims against case law and federal statutes, and apply the Court’s jurisprudence. Finally, they prepare a document called Instrução (Portuguese for Instruction) that summarizes the case, the evidence, and the legal analysis and proposes a decision. The instruction is then sent to the court’s judicial panel for final adjudication. Despite the variability in human effort required for each case, it can be reasonably estimated that the review of a single case necessitates approximately 30 hours, incurring a cost of approximately \$1,750 before the case is even presented to the court’s judicial panel. Given an annual average processing rate of about 2,000 cases within the purview of the INACIA project and the textual nature of the task, audit case processing is an ideal candidate for an LLM application that would help expedite review while reducing costs. [^6] For comparison purposes, INACIA processes a case in under 20 minutes and at an initial cost of \$10, subject to further human review. A mere 1 percent reduction in the time required for a human to process a case would suffice to offset the costs associated with INACIA.
@@ 3.1. Overview
Figure 1 summarizes INACIA’s main case processing pipeline, showcasing the holistic nature of case analysis. While there is a direct, sequential flow (i.e., solid arrows) of tasks (i.e., blue boxes) needed to produce the instruction (case suggested decision), the dashed lines show the association (or dependency) between tasks. The lighter grey boxes indicate data objects needed to start processing or to be produced as intermediate or final products for other TCU systems. In the next subsections, we describe each element of this pipeline in detail.
Figure 1. INACIA pipeline.
@@ 3.2. Inputs
The pipeline starts off ingesting all documents available at case submission (“Case documents” grey box). There are two sets of documents in the submission package: (1) the main document is the representação (in Portuguese), which is in nature similar to a civil complaint. It is a formal document laying out the claims of wrongdoing and why TCU should take the case; and (2) supporting documents containing the evidence of wrongdoing as presented by the plaintiff (e.g., if the plaintiff alleges that the defendant has not complied with a contract, the contract is a document that supports the case).
@@
Both the representação and supporting documents are stored in multiple formats, and they can be classified in one of two categories: (1) searchable files:electronically-created files with searchable content (e.g., native PDF document or Word document printed as PDF); (2) unsearchable files:scanned physical files or images whose content is not searchable (e.g., a photo of a contract or signature). Since INACIA uses the textual content of case documents to perform its tasks, we needed to implement two data extraction solutions to handle all files stored by TCU. These solutions needed to deliver high-quality extraction at a low cost for the solution to be sustainable once TCU takes the system to production.
@@
Most files in TCU’s case database are searchable and primarily stored as PDF documents. Thus, the first implemented solution is a standard PDF extraction pipeline using the Apache Tika framework (Apache, [n. d.]), which can handle all types of searchable files in TCU’s case database. Besides being an open-source tool (thus meeting the low-cost criterion), Tika delivered better results in text extraction than PyPDF2 (Fenniak, [n. d.]), another popular framework for processing searchable PDF documents. We compared and contrasted Tika and PyPDF2 in the following ways: (1) which framework was better at extracting textual content and its structure (e.g., a two-column document processed as separate, two columns instead of a single column with left-to-right text mixing up the two columns). (2) which framework was better at extracting characters from different encoding formats (e.g., those not present in UTF-8, Latin-1, or Windows-1252 encodings).
Tika performed better in both criteria – we particularly liked that it reports the count of invalid characters in the processed text as extraction metadata, allowing us to create a measure of extraction quality by easily computing the percentage of invalid characters over all characters extracted.
@@
Whenever Tika fails in text extraction (e.g., it reported zero searchable characters after processing the text), INACIA then deploys the second text extraction solution, which relies on Microsoft’s Azure Form Recognizer service for Optical Character Recognition (OCR) (Microsoft, [n. d.]). Form Recognizer will ingest any unsearchable file and return the extracted text in plain format. We also compared Form Recognizer to other industry-leading OCR frameworks (Tesseract (Smith, 2007), and Google Cloud Vision (Google, [n. d.])) using the high-quality and low-cost criteria. While they are similarly priced, Form Recognizer had higher-quality results and richer extraction metadata (e.g., information on the position and content of tables).777 To assess text extraction from either document type, we evaluated results qualitatively. We manually compared extracted elements from case documents, such as the position of tables, columns, and text, allowing for a more nuanced understanding of the strengths and weaknesses of each tool in practical scenarios. We evaluated a sample of documents from TCU’s database containing all cases and their documents between 2019 and 2022. We pulled out a minimum of five documents per year. Within these documents, we had at least one document of easy, medium, and hard extraction, determined by the number of pages and their content as follows: (1) Easy: up to 10-page documents with minimal tables; (2) Medium: up to 25-page documents with structured data or images; (3) Hard: over 25-page documents with structured and unstructured data elements, images, or handwritten content.
@@ 3.3.1. Basic Information Extraction
Figure 2. Approach used for Basic Information Extraction.
Figure 2 illustrates the basic information extraction process. It is a structured and systematic process relying on prompt engineering and retrieval-augmented generation (RAG) techniques (Lewis et al., 2020): we will first feed the LLM a structured list of the information we need to extract from case documents (“Pre-defined form” green box). This list was constructed using older cases and their representação containing the information TCU needs to retain for record-keeping purposes 888 “Basic information” examples are the case ID, case type, plaintiff name, plaintiff ID, etc. . We also send over detailed explanations of what each of the variables is. Finally, we append to our combined prompt from above the first two pages of a case’s representação (“Main document” gray box), which should contain most of the information needed.
We deploy our RAG workflow when we cannot find all necessary information in the first pages of the representação. Along with extracting all textual content from case documents, we created searchable indexes of case documents. We allowed the LLM to reason through the form and query such indexes if it did not find the information in the first two pages of the main case document. An example that helps illustrate this mechanism is that TCU staff often need to look at federal agencies’ procurement documents to evaluate whether they have followed all necessary rules when contracting external vendors. INACIA can search supporting documents of a case to find procurement contracts and retrieve the necessary information (the rightmost side of Figure 2 illustrates this process). We use standard industry-leading search algorithms, such as BM25 (Robertson et al., 1994), to retrieve the relevant passages from the indexed case documents. The LLM then processes these passages to extract answers, ensuring a streamlined and consistent approach. The final step is retrieving the remaining information using BM25 and manually crafted queries. The resulting information is then fed into the LLM for the final extraction, completing the form. In this workflow, the LLM functions not as an intelligent agent but rather as a tool within the pipeline for interpreting search results.

### 26 Feb 2024 v3

@@ Abstract.
This paper introduces INACIA (Instrução Assistida com Inteligência Artificial), a groundbreaking system designed to integrate Large Language Models (LLMs) into the operational framework of Brazilian Federal Court of Accounts (TCU). The system automates various stages of case analysis, including basic information extraction, admissibility examination, Periculum in mora and Fumus boni iuris analyses, and recommendations generation. Through a series of experiments, we demonstrate INACIA’s potential in extracting relevant information from case documents, evaluating its legal plausibility, and formulating propositions for judicial decision-making. Utilizing a validation dataset alongside LLMs, our evaluation methodology presents a novel approach to assessing system performance, correlating highly with human judgment. These results underscore INACIA’s potential in complex legal task handling while also acknowledging the current limitations. This study discusses possible improvements and the broader implications of applying AI in legal contexts, suggesting that INACIA represents a significant step towards integrating AI in legal systems globally, albeit with cautious optimism grounded in the empirical findings.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: DGOVjournalvolume: 37journalnumber: 4article: 111publicationmonth: 2ccs: Applied computing Lawccs: Information systems Information extractionccs: Computing methodologies Natural language processingccs: General and reference Empirical studies
@@ 1. Introduction
In this paper, we introduce a groundbreaking application of LLMs for administrative case processing and assisted decision-making developed jointly by the Brazilian Federal Court of Accounts (TCU, or Tribunal de Contas da União, in Portuguese) and experts at several Brazilian research organizations. INACIA (Instrução Assistida com Inteligência Artificial, in Portuguese, or AI-Assisted Legal Instruction) is a system that suggests adjudication directions to court decision-makers – consisting of structured reasoning around cases’ claims, evidence, and relationship to legal provisions in Brazilian law. TCU’s staff auditors draw from INACIA’s analytical support but ultimately bear the responsibility of making adjudication recommendations to court decision-makers. Its inputs are case filings; outputs are suggestions to the auditor to write the initial instruction [^1] The initial instruction is the document that contains the directions written by the auditor, including the main facts of the case, the legal basis, and the adjudication proposals.. INACIA uses a combination of pre-trained large language models (LLMs) and search engines to extract information from case documents and to reason through the extracted information.
Importantly, INACIA goal’s is to expedite case adjudication as it sifts through case documents and connects claims, evidence, and legal provisions. INACIA, however, is not replacing skilled auditors; instead, it is augmenting their productivity. By handling routine and straightforward tasks, INACIA allows auditors to focus on more complex, analytical tasks, thereby deepening the overall effectiveness and efficiency of TCU staff.
@@
The remainder of this paper is organized as follows. Section 2 provides an overview of TCU and the role of AI in its operations. Section 3 reviews the related work in the field of AI and LLMs in legal settings. Section 4 introduces INACIA, detailing its case processing pipeline and the challenges it addresses. Section 5 presents the experiments conducted to evaluate INACIA’s performance, including the validation dataset and the evaluation methodology. Section 6 discusses the results of the experiments and their implications. Section 7 discusses the broader implications of INACIA and the challenges of integrating AI in legal systems. Finally, Section 8 concludes the paper and outlines future research directions.
2. Application Domain
@@ 2.3. Audit Case Processing
TCU investigates (and eventually issues rulings on) the use of federal funds in two ways: (1) when TCU independently audits federal agencies (or anyone using federal funds) or (2) when TCU receives complaints filed by individuals, private entities, or public agencies on the misuse of federal funds.
@@
While case proceedings may vary based on the initiation method, case processing is typically a time-intensive task primarily handled by TCU staff. Auditors must review case filings, confirm whether TCU should hear the case, search and append external evidence of wrongdoing (if necessary), check claims against case law and federal statutes, and apply the Court’s jurisprudence. Finally, they prepare a document called Instrução (Portuguese for Instruction) that summarizes the case, the evidence, and the legal analysis and proposes adjudication recommendations. The instruction is then sent to the court’s judicial panel for final adjudication. Despite the variability in human effort required for each case, it can be reasonably estimated that the review of a single case necessitates approximately 30 hours, incurring a cost of approximately \$1,750 before the case is even presented to the court’s judicial panel. Given an annual average processing rate of about 2,000 cases within the INACIA project’s purview and the task’s textual nature, audit case processing is an ideal candidate for an LLM application that would help expedite review while reducing costs. [^3] For comparison purposes, INACIA processes a case in under 20 minutes and at an initial cost of \$10, subject to further human review. A mere 1 percent reduction in the time required for a human to process a case would suffice to offset the costs associated with INACIA.
2.4. Large Language Models in Brazilian Legal Settings
Though natural language processing (NLP) applications have been familiar to courts both in Brazil and worldwide for some time, the use of large language models as the engine behind NLP applications is relatively new – only recently courts across the world have begun using LLMs for document summarization, chatbots, writing briefs, etc. (Rabihah Butler, 2024)
TCU is a particularly interesting use case as it predates the popularization of LLMs through OpenAI GPT models in late 2022 and throughout 2023. Acknowledging the rapid evolution of AI and in line with its mission to foster innovation and enhance public administration, TCU carried out a series of pre-commercial procurement (PCP) studies between 2019 and 2022 to design a public procurement call for an AI-powered solution to expedite audit case processing. Open for bidding in early 2022, this procurement call was the first of its kind in Brazil as it sought to contract a pre-market AI product requiring research and development and carrying significant technological risk (the solution envisioned by TCU should go beyond the state-of-the-art NLP tools available then).
INACIA was the winning proposal of this procurement call. A team of researchers at the State University of Campinas (UNICAMP) and start-ups NeuralMind and Terranova Consultoria put together the only proposal to use an LLM to handle all NLP tasks needed in the audit processing product (detailed in the following sections). Back in early 2022, LLMs were fairly unknown outside of the NLP expert community and were not consensually seen as the way forward in NLP research. TCU, however, trusted the INACIA proposal, and this manuscript is based on the work done during the proof of concept (POC) development stage.
TCU goals are to focus auditors’ expertise on the most challenging tasks of case processing while handing over the more routine tasks to the AI system. This strategy is well-aligned with the evidence on how best to incorporate technology for productivity gains and labor output augmentation (Acemoglu and Autor, 2011). The current expectation is that AI will change how TCU handles cases from a sequential, step-by-step approach to a more holistic approach in which filings are evaluated simultaneously and connected to external resources necessary for decision-making from the outset. Effectively, this system would eliminate redundancies and streamline case processing without requiring regulation change.
3. Related Work
Artificial Intelligence (AI) and Large Language Models (LLMs) are now part of products and services across various jurisdictions and domains, highlighting a global movement towards technologically enhanced legal systems. Some of these include Chalkidis et al. (2020), who developed Legal-BERT, a domain-specific adaptation of BERT for legal documents, demonstrating the effectiveness of language models in legal document analysis and information extraction. Similarly, Safaei and Longo (2023) focused on zero-shot text matching for financial auditing using LLMs, showcasing the potential of AI in improving the accuracy and efficiency of legal and financial document processing. The work of Hillebrand et al. (2023) further extends this by presenting ZeroShotALI, a system that leverages LLMs for recommending relevant legal texts, underlining the advancements in AI-driven legal recommendation systems.
Additionally, the exploration of GPT models in zero-shot settings for legal text annotation by Savelka (2023) provides insights into the capabilities of LLMs in the semantic understanding and classification of legal documents. Paul et al. (2023) complements this investigation into pre-training LLMs on Indian legal corpus, which enriches the models’ understanding of legal texts from different legal systems and highlights the adaptability and scalability of AI techniques in diverse legal environments.
Finally, Sipra et al. (2024) assesses the challenges and opportunities in AI-driven criminal justice systems, addressing ethical, transparency, and accountability issues, which are critical in the application of AI within legal systems. These works collectively underscore the transformative potential of AI and LLMs in legal research, practice, and administration, offering valuable insights for the integration of such technologies into the Brazilian Federal Court of Accounts (TCU) through INACIA.
@@ 4.1. Overview
As stated previously, INACIA’s goal is to assist TCU auditors in streamlining case adjudication while refocusing their tasks to high-value-added work, e.g., replacing routine tasks of information extraction for tasks requiring critical reasoning over the combination of a case’s claims, the applicable legal framework and evidence brought forth by plaintiffs. The main advantage of using pre-trained LLMs is that they can work in a few-shot or zero-shot setting, meaning that they can be used to process cases without the need for training on a specific dataset or domain. This is particularly important in the TCU context, as the number of cases is relatively small compared to the datasets usually used to train LLMs. OpenAI’s models, for example, are trained on a huge dataset of internet text, [^4] OpenAI does not disclose the exact size of the dataset used to train its models, but it is known to be in the order of hundreds of gigabytes. which allows them to understand and generate text in a wide range of domains and languages. Using search engines, on the other hand, allows INACIA to access external information, such as case law and legal provisions, which are not present in the case documents and potentially not present (or present in an insignificant amount) in the training data of the LLMs. It allows INACIA to overcome one of the most common limitations of LLMs, which is the lack of access to external information and the tendency to generate text that is not grounded in reality (a.k.a. hallucination).
Figure 1 depicts INACIA’s main case processing pipeline, showcasing the holistic nature of case analysis and how much of the work done by auditors can be replaced by INACIA work. While there is a direct, sequential flow (i.e., solid arrows) of tasks (i.e., blue boxes) needed to produce the instruction (case adjudication suggestion), the dashed lines show the association (or dependency) between tasks. The lighter grey boxes indicate data objects needed to start processing or to be produced as intermediate or final products for other TCU systems.555 INACIA is available to TCU in two formats: (1) a Python API that can be executed end-to-end from an interactive shell, e.g., a Jupyter Notebook, or from source or via classes and methods that handle each component of the pipeline as an independent module; and (2) a staging environment developed for presenting INACIA’s proof of concept to the business team at TCU. In either case, the outputs of one INACIA module are the inputs to the next module. INACIA is currently designed to run end-to-end with no human intervention – though its execution could be broken out and interspersed with manual checkpoints along the way. The backend of INACIA is entirely structured in the Azure Cloud and the user has the option of choosing any of the pre-trained language models available by Microsoft on its Azure Portal. In the next subsections, we describe each element of this pipeline in detail.
Figure 1. End-to-end INACIA pipeline (from case documents to initial instruction).
@@ 4.2. Inputs
The pipeline starts off ingesting all documents available at case submission (“Case documents” grey box). There are two sets of documents in the submission package: (1) the main document is the representação (in Portuguese), which is in nature similar to a civil complaint. It is a formal document laying out the claims of wrongdoing and why TCU should take the case; and (2) supporting documents containing the evidence of wrongdoing as presented by the plaintiff. For example, if the plaintiff alleges that the defendant has not complied with a contract, the contract is a document that supports the case.
None of these documents follow a strict format. Lawyers are free to petition TCU in any way as long as the main sections of their complaint follow the guidelines set by the TCU. The representação, thus, is a free-form text document and the supporting documents can be anything (from pictures to invoices to full procurement contracts). There is no set range of pages for either type of document. Figure 2 contains snapshots of files typically found in TCU’s case database. Figure 1(a) shows a representação in a native PDF format, and Figure 1(b) shows a supporting document (a business incorporation document) in a scanned image format.
(a) A representação
(b) A supporting document (business incorporation documents)
Figure 2. Examples of files typically found in TCU’s case database.
@@
Both the representação and supporting documents are stored in multiple formats, and they can be classified in one of two categories: (1) searchable files:electronically-created files with searchable content (e.g., native PDF document or MS Word document printed as PDF); (2) unsearchable files:scanned physical files or images whose content is not searchable (e.g., a photo of a contract or signature). Since INACIA uses the textual content of case documents to perform its tasks, we needed to implement data extraction solutions to handle all file types. These solutions needed to deliver high-quality extraction at a low cost for the solution to be sustainable once TCU takes the system to production.
@@
Most files in TCU’s case database are searchable and primarily stored as PDF documents. Thus, the first implemented solution is a standard PDF extraction pipeline using the Apache Tika framework (Apache, [n. d.]), which can handle all types of searchable files in TCU’s case database. Besides being an open-source tool (thus meeting the low-cost criterion), Tika delivered better results in text extraction than PyPDF2 (Fenniak, [n. d.]), another popular framework for processing searchable PDF documents. We compared and contrasted Tika and PyPDF2 in the following ways: (1) which framework was better at extracting textual content and its structure (e.g., a two-column document processed as separate, two columns instead of a single column with left-to-right text mixing up the two columns). (2) which framework was better at extracting characters from different encoding formats (e.g., those not present in UTF-8, Latin-1, or Windows-1252 encodings). Tika performed better in both criteria – we particularly liked that it reports the count of invalid characters in the processed text as extraction metadata, allowing us to create a measure of extraction quality by computing the percentage of invalid characters over all characters extracted.
@@
Whenever Tika fails in text extraction (i.e., it reported zero searchable characters after processing the text), INACIA then deploys the second text extraction solution, which relies on Microsoft’s Azure Form Recognizer service for Optical Character Recognition (OCR) (Microsoft, [n. d.]). Form Recognizer ingests any unsearchable file and returns the extracted text in plain format. We also compared Form Recognizer to other industry-leading OCR frameworks (Tesseract (Smith, 2007), and Google Cloud Vision (Google, [n. d.])) using the high-quality and low-cost criteria. While they are similarly priced, Form Recognizer had higher-quality results and richer extraction metadata (e.g., information on the position and content of tables). To assess text extraction from either document type, we evaluated results qualitatively. We manually compared extracted elements from case documents, such as the position of tables, columns, and text, allowing for a more nuanced understanding of the strengths and weaknesses of each tool in practical scenarios. We evaluated a sample of documents from TCU’s database containing all cases and their documents between 2019 and 2022. We pulled out a minimum of five documents per year. Within these documents, we had at least one document of easy, medium, and hard extraction, determined by the number of pages and their content as follows: (1) Easy: up to 10-page documents with minimal tables; (2) Medium: up to 25-page documents with structured data or images; and (3) Hard: over 25-page documents with structured and unstructured data elements, images, or handwritten content.
@@ 4.3.1. Basic Information Extraction
Figure 3. Pipeline used for Basic Information Extraction.
Figure 3 illustrates the approach used for basic information extraction. The solid lines indicate the flow of information between components of the system (in grey are the data objects, in blue the agent components; in orange technologies external to the main INACIA pipeline); the work done by INACIA’s basic information extraction agent is depicted by the bounded box. It is a structured and systematic process relying on prompt engineering and retrieval-augmented generation (RAG) techniques (Lewis et al., 2020). First, we feed the LLM a structured list of the information we need to extract from case documents (“Pre-defined form” green box). This list was constructed using older cases and their representação containing the information TCU needs to retain for record-keeping purposes.666 “Basic information” examples are the case ID, case type, plaintiff name, plaintiff ID, etc. We also provide the model with detailed explanations of what each of the variables is. Finally, we append to our combined prompt from above the first two pages of a case’s representação (“Main document” gray box), which should contain most of the information needed.
We deploy our RAG workflow when we cannot find all necessary information in the first pages of the representação. Along with extracting all textual content from case documents, we created searchable indexes of case documents. We allowed the LLM to reason through the form and query such indexes if it did not find the information in the first two pages of the main case document. An example that helps illustrate this mechanism is that TCU staff often need to look at federal agencies’ procurement documents to evaluate whether the agencies have followed all necessary rules when contracting external vendors. INACIA can search supporting documents of a case to find procurement contracts and retrieve the necessary information (the rightmost side of Figure 3 illustrates this process). We use standard industry-leading search algorithms, such as BM25 (Robertson et al., 1994), to retrieve the relevant passages from the indexed case documents. The LLM then processes these passages to extract answers, ensuring a streamlined and consistent approach. For retrieving the remaining information using BM25, we manually crafted pre-defined queries. The resulting information is then fed into the LLM for the final extraction, completing the form. In this workflow, the LLM functions not as an intelligent agent but rather as a tool within the pipeline for interpreting search results.